OpenStudy (anonymous):

If $2000 is interest at an interest rate of 4.5% per year, compounded daily, find the value investment after a given number of years. (round to the nearest decimal)

OpenStudy (tkhunny):

Annual \(2000\cdot(1 + 0.045)^{n}\) Semi-Annual \(2000\cdot\left(1 + \dfrac{0.045}{2}\right)^{n*2}\) Quarterly \(2000\cdot\left(1 + \dfrac{0.045}{4}\right)^{n*4}\) Monthly \(2000\cdot\left(1 + \dfrac{0.045}{12}\right)^{n*12}\) Weekly \(2000\cdot\left(1 + \dfrac{0.045}{52}\right)^{n*52}\) Are you seeing a theme?

OpenStudy (anonymous):

would daily be 365?

OpenStudy (tkhunny):

Yes.